OpenStudy (tester97):

help on 3 questions?

OpenStudy (skyz):

(¬‿¬)

OpenStudy (tester97):

here are the 3 questions

OpenStudy (tester97):

idk from my question?

OpenStudy (anonymous):

Why do you ask such simple questions :>

OpenStudy (tester97):

can u help me peter?

OpenStudy (skyz):

Waka waka waka waka waka waka waka

OpenStudy (anonymous):

this time for... anyway, sure, why not? ^.^

OpenStudy (anonymous):

(5 - 5i) - (-6 - 3i) This is just plain algebra isn't it? Add and subtract like terms and all that?

OpenStudy (tester97):

wat?

OpenStudy (anonymous):

wat?

OpenStudy (anonymous):

let's pick another letter, maybe that 'i' has you thrown off or something... I choose p :D What if it was (5 - 5p) - (-6 - 3p) instead, would it be a problem?

OpenStudy (tester97):

no thats pretty easy actually

OpenStudy (anonymous):

Well then, show me, what's the answer if it were p instead?

OpenStudy (tester97):

just foil right?

OpenStudy (anonymous):

To heck with foil, this is just subtraction! (LOL and not multiplication ^.^ )

OpenStudy (tester97):

sorry im so lost its been a long day...........just subtract the 5p from 5? and the 3p from 6?

OpenStudy (anonymous):

You can't subtract unlike terms, LOL

OpenStudy (tester97):

well can u explain in further detail what i am supposed to do?

OpenStudy (anonymous):

Okay... how about an example? ^.^

OpenStudy (anonymous):

(4 + 3i) - (-7 + i)

OpenStudy (tester97):

ok

OpenStudy (anonymous):

First, get rid of the parentheses... the ones on the left, you can just do away with. 4 + 3i - (-7 + i) Now, since it's a minus-sign in the middle, when you remove the parentheses on the right-side, you change the sign of everything inside: 4 + 3i + 7 - i Now just combine like terms, the 4 and 7, as well as the 3i and -i = 11 + 2i get it?

OpenStudy (tester97):

yes so far.......go on.........

OpenStudy (anonymous):

That's it. You just apply the same principle to your question.

OpenStudy (tester97):

wait so thats the answer that simple?

OpenStudy (anonymous):

that's the answer to my example. answering your question requires a similar process ^.^

OpenStudy (tester97):

ok let me solve those questions will u tell me if they are correct?

OpenStudy (anonymous):

sure.

OpenStudy (tester97):

ok for #3 i got 11-2i right or wrong?

OpenStudy (anonymous):

It's correct ^.^

OpenStudy (tester97):

ok but i tried to work out #4 but i dont get it can u help?

OpenStudy (anonymous):

THAT's where you use FOIL.

OpenStudy (tester97):

ohhhhhh

OpenStudy (anonymous):

Because it's multiplication. Do remember that after the dust settles with your FOIL that \[\Large i^2 = -1\]

OpenStudy (tester97):

sooooo now i have to multiply is by neg 1?

OpenStudy (anonymous):

worry about that later, for now, just FOIL away.

OpenStudy (tester97):

i got \[-12i ^{2}-12i+16\]

OpenStudy (tester97):

now what?

OpenStudy (anonymous):

a bit off... could you FOIL again? only this time, more carefully?

OpenStudy (tester97):

-4-4i?

OpenStudy (anonymous):

that's strange, where did the i^2 go? Don't skip steps LOL That's how we go wrong :D

OpenStudy (tester97):

idk im so confused

OpenStudy (tester97):

i wll try again

OpenStudy (tester97):

-12i^2-4i+8?

OpenStudy (anonymous):

That's good :) NOW use the fact that \(\large i^2 = -1\) and simplify further ^.^

OpenStudy (tester97):

so multiply -12 times -1?

OpenStudy (anonymous):

You got that right ^.^

OpenStudy (anonymous):

You got your answer?

OpenStudy (tester97):

12-4i+8?

OpenStudy (anonymous):

yes... go on... simplify the darn thing :>

OpenStudy (tester97):

simplifies to 20-4i

OpenStudy (anonymous):

Now was that so hard? ;)

OpenStudy (tester97):

OHMEHFREAKINGGOOSENEEKS U ARE AMAZING PETER

OpenStudy (anonymous):

Tell me something I don't know :)

OpenStudy (tester97):

ok on the last one i have to divide it? or what?

OpenStudy (anonymous):

oh, division... this is the tricky one... but not that tricky. ever heard of conjugates? :)

OpenStudy (tester97):

conjuwhats?

OpenStudy (anonymous):

Conjugate of a complex number. It sounds... complex, but it really isn't. It just means you change the sign of the imaginary part. For example: conjugate of 1+3i is 1-3i conjugate of 4 - i is 4+i conjugate of 5i is -5i conjugate of 7 is 7 (since there's no imaginary part) In general, conjugate of a + bi is a - bi get it?

OpenStudy (tester97):

so like the opposite of the first thing is the second?

OpenStudy (anonymous):

uhh... don't confuse me with words like opposite... stick to conjugate in this context. Here's a little drill...find the conjugates of these numbers: 1-2i 5+9i 7-i

OpenStudy (tester97):

1+2i 5-9I 7+1 ?/???

OpenStudy (anonymous):

Correct. See? Not very complex at all... ready to work on your third question now?

OpenStudy (anonymous):

except for a few typos... it should be 7+i, not 7+1 -_-

OpenStudy (tester97):

sorry but i think i got it other than those pesky typos >:(

OpenStudy (anonymous):

Okay, when dividing complex numbers. we don't want any i in the denominator. So what we do is \[\Large \frac{-6+i}{-5+i}\]

OpenStudy (anonymous):

We multiply BOTH the numerator and denominator by the CONJUGATE of the denominator. What's the conjugate of the denominator?

OpenStudy (tester97):

-5+i?

OpenStudy (anonymous):

that's the denominator itself... but what's its CONJUGATE? c'mon, you can do this ^.^

OpenStudy (tester97):

whoops i meant to say -5-i ;)

OpenStudy (anonymous):

That's good. Now multiply both the numerator and denominator by that conjugate: \[\Large \frac{-6+i}{-5+i}\times \frac{-5-i}{-5-i}\] FOIL twice. Just a heads' up: When you multiply a complex number by its own conjugate, you SHOULD get A REAL NUMBER. okay? good. Proceed...

OpenStudy (anonymous):

Here, I'll make it simpler for you :) \[\Large \frac{(-6+i)(-5-i)}{(-5+i)(-5-i)}\]

OpenStudy (tester97):

give meh a minute............

OpenStudy (anonymous):

On the denominator, it might help you to remember that \[\large (x+y)(x-y) = x^2-y^2\]

OpenStudy (tester97):

\[\frac{ 30+i }{ 25-i^2 }\] is this right?

OpenStudy (anonymous):

Denominator is correct... numerator... is not :D By the way, you can further simplify the denominator... remember \[\large i^2 = -1\]

OpenStudy (tester97):

let me try agian

OpenStudy (anonymous):

Please do ^.^

OpenStudy (tester97):

-i^2-i+30?

OpenStudy (anonymous):

Still no... please be very careful with signs :) And please include your denominator too :)

OpenStudy (tester97):

ok im getting the same answer of 30+i-1^2/25-i^2 what am i doing wrong? :(

OpenStudy (anonymous):

because your earlier numerator had a -i :P

OpenStudy (anonymous):

is that 1^2?

OpenStudy (tester97):

sorry i mean i^2 im still getting used to using the i

OpenStudy (anonymous):

Okay, now it's correct :) \[\Large \frac{30+i-i^2}{25-i^2}\] Now, simplify... and remember \[\large i^2=-1\] finish it!

OpenStudy (tester97):

ok i can do this!

OpenStudy (skyz):

ALLEY

OpenStudy (skyz):

ISNT NICOLE PRETTY BEAUTIFUL C:

OpenStudy (tester97):

shhhhhhh im tryna learn

OpenStudy (anonymous):

hah Pan 1, Sky 0 :>

OpenStudy (skyz):

The Sky will fall upon you Peter :c

OpenStudy (anonymous):

But today is not that day... @tester97 got your answer yet? :D

OpenStudy (tester97):

is it \[\frac{ 31+i }{ 26}\]

OpenStudy (anonymous):

hurrah :D

OpenStudy (anonymous):

That wasn't so bad, was it? ^.^

OpenStudy (tester97):

:OIM FINALLY DONE THANKS SOOOOOOOOOOOOOOOO MUCH PETER!

OpenStudy (anonymous):

No problem ^.^